The PDE 2∂

2

z

∂x

2


+2∂

2

z

∂y

2


+4∂

2

z

∂x∂y


=2x

2∂2z∂x2+2∂2z∂y2+4∂2z∂x∂y=2x,

is 


1
Expert's answer
2021-12-09T17:25:32-0500

The PDE 


"2\frac{∂^2z}{∂x^2}+2\frac{∂^2z}{∂y^2}+4\frac{∂^2z}{∂x∂y}=2x^2\frac{∂^2z}{∂x^2}+2\frac{∂^2z}{∂y^2}+4\frac{∂^2z}{∂x∂y}=2x"


second order pde can be written in the form:

"az_{xx}+2bz_{xy}+cz_{yy}+dz_x+ez_y+fz=g(x,y)"


for "2\frac{∂^2z}{∂x^2}+2\frac{∂^2z}{∂y^2}+4\frac{∂^2z}{∂x∂y}=2x" :


"b^2-ac=4-4=0"

this is parabolic equation



for "2x^2\frac{∂^2z}{∂x^2}+2\frac{∂^2z}{∂y^2}+4\frac{∂^2z}{∂x∂y}=2x" :


"b^2-ac=4-4x^2"


then:

for "x=0" :

this is parabolic equation


for "x<-2" and "x>2" :

"b^2-ac<0"

this is elliptic equation


for "-2<x<2" :

"b^2-ac>0"

this is hyperbolic equation